Unknown concentration of 100.0 mL diprotic acid is titrated against 1.0 M NaOH. The initial burette reading was 1.3 ml, and the final burette reading during the end point was 19.6 ml. What is the concentration of the acid?



1
Expert's answer
2022-01-18T08:11:02-0500

Let Concentration of Diprotic Acid is Y M .

100 ml of Y M Diprotic Acid have

Milli Moles of Diprotic Acid = 100 × Y .

Milli geq of Diprotic Acid = 100 × Y × 2

= 200 Y .

Volume used for NaOH = ( 19.6 - 1.3 ) ml .

Molarity of NaOH = 1 M .

Millimoles of NaOH = ( 18.3 × 1 ) .millimoles

Milli geq = 18.3


Milligeq of Acid and Milligeq of base should be equal .


200 Y = 18.3


Y = 18.3 / 200 = 0.0915 M .